# MMMU-Pro: A More Robust Multi-discipline Multimodal Understanding Benchmark

### v1

@@ 4 Related Work
Multimodal Large Language Models. Recent progress in multimodal AI has been marked by innovative training approaches (Lu et al., 2019; Tan and Bansal, 2019; Chen et al., 2020; Zhou et al., 2020; Zhang et al., 2021; Li et al., 2020; Alayrac et al., 2022; Awadalla et al., 2023). Inspired by the success of large language models, researchers have developed various models with improved instruction-following capabilities (Liu et al., 2023c, b, 2024a; Li et al., 2024b; Dai et al., 2023; Zhu et al., 2023; Zhang et al., 2023; Gao et al., 2023; Ye et al., 2023a, b; Zhao et al., 2023; Li et al., 2023; Monajatipoor et al., 2023; Zhao et al., 2024; Li et al., 2024c; Lin et al., 2024; Zhang et al., 2024). Proprietary models such as GPT-4V (OpenAI, 2023), GPT-4o  (OpenAI, 2024a), Gemini  (Team et al., 2023), and Claude-3.5  (Anthropic, 2024) have demonstrated strong performance across various vision-language tasks. However, a significant challenge remains in accurately evaluating the capabilities of these advanced multimodal models, highlighting the need for more robust and comprehensive benchmarks.
MLLM Benchmarks. The rise of more advanced multimodal pre-training and instruction tuning has exposed the limitations of earlier benchmarks like VQA (Antol et al., 2015; Goyal et al., 2017), OK-VQA (Marino et al., 2019), and MSCOCO (Lin et al., 2014), which no longer suffice to evaluate the full spectrum of LMMs capabilities. To address this, recent benchmarks such as LAMM (Yin et al., 2023b), LVLM-eHub (Xu et al., 2023), SEED (Li et al., 2024a), MMBench (Liu et al., 2023d),CV-Bench (Tong et al., 2024a), MM-Vet (Yu et al., 2024), Mantis (Jiang et al., 2024), and BLINK (Fu et al., 2024) have emerged, covering aspects from basic perception to hallucination detection (Cui et al., 2023; Liu et al., 2023a). However, existing benchmarks often fall short in evaluating expert-level domain knowledge and complex reasoning. While MMMU (Yue et al., 2024) made strides by incorporating multimodal, college-level questions, it still permits text-only models to find shortcuts (Lu et al., 2023). To overcome these limitations, we introduce MMMU-Pro, a more robust version specifically designed to better evaluate multimodal reasoning by eliminating text-only answerable questions, expanding candidate options, and introducing a vision-only input setting that mirrors real-world scenarios where text and images are naturally intertwined.

### v2

@@ 4 Related Work
Multimodal Large Language Models. Recent progress in multimodal AI has been marked by innovative training approaches (Lu et al., 2019; Tan and Bansal, 2019; Chen et al., 2020; Zhou et al., 2020; Zhang et al., 2021; Li et al., 2020; Alayrac et al., 2022; Awadalla et al., 2023). Inspired by the success of large language models, researchers have developed various models with improved instruction-following capabilities (Liu et al., 2023c, b, 2024a; Li et al., 2024b; Dai et al., 2023; Zhu et al., 2023; Zhang et al., 2023; Gao et al., 2023; Ye et al., 2023a, b; Zhao et al., 2023; Li et al., 2023; Monajatipoor et al., 2023; Zhao et al., 2024; Li et al., 2024c; Lin et al., 2024; Zhang et al., 2024a). Proprietary models such as GPT-4V (OpenAI, 2023), GPT-4o  (OpenAI, 2024a), Gemini  (Team et al., 2023), and Claude-3.5  (Anthropic, 2024) have demonstrated strong performance across various vision-language tasks. However, a significant challenge remains in accurately evaluating the capabilities of these advanced multimodal models, highlighting the need for more robust and comprehensive benchmarks.
MLLM Benchmarks. The rise of more advanced multimodal pre-training and instruction tuning has exposed the limitations of earlier benchmarks like VQA (Antol et al., 2015; Goyal et al., 2017), OK-VQA (Marino et al., 2019), and MSCOCO (Lin et al., 2014), which no longer suffice to evaluate the full spectrum of LMMs capabilities. To address this, recent benchmarks such as LAMM (Yin et al., 2023b), LVLM-eHub (Xu et al., 2023), SEED (Li et al., 2024a), MMBench (Liu et al., 2023d),CV-Bench (Tong et al., 2024a), MM-Vet (Yu et al., 2024), Mantis (Jiang et al., 2024), and BLINK (Fu et al., 2024) have emerged, covering aspects from basic perception to hallucination detection (Cui et al., 2023; Liu et al., 2023a). However, existing benchmarks often fall short in evaluating expert-level domain knowledge and complex reasoning (Lu et al., 2023a; Zhang et al., 2024b). While MMMU (Yue et al., 2024) made strides by incorporating multimodal, college-level questions, it still permits text-only models to find shortcuts (Lu et al., 2023b; Zhang et al., 2024b). To overcome these limitations, we introduce MMMU-Pro, a more robust version specifically designed to better evaluate multimodal reasoning by eliminating text-only answerable questions, expanding candidate options, and introducing a vision-only input setting that mirrors real-world scenarios where text and images are naturally intertwined.